OpenStudy (anonymous):

Which of these is NOT the equivalent to 13/15 ? 54/60 143/165 117/135 26/30

OpenStudy (anonymous):

the first one because if u solve the others u'll the same decimal .86667

OpenStudy (callisto):

Once again, simplify all the numbers there For example, 13/15 cannot be simplified anymore 54/60 = (6x9)/(6x10) = 9/10 Can you do the rest?